WHAT GRAM OF SODIUM IS PRESENT IN 11.7g OF SODIUM CHLORIDE?


 

Grade:9

1 Answers

vikas askiitian expert
13 years ago

1mol NaCl gives 1molCl &  1molNa

 1mole NaCl =(23+35.5)gmNaCl

                   =58.5gmNaCl

that means 58.5gm NaCl gives 1 mol of Na=23gm

          1gmNaCl=23/58.5 gm Na

       11.7gmNaCl=23x11.7/58.5=4.6gm Na           ans